I would just like to verify this is correct. My answer only leaves 6 degrees for the last angle, which doesn't seem right.
Thanks!
The diagonal brace on a rectangular gate forms an angle with the horizontal side with degree x and an angle with the vertical side with degree measure 2x – 3. Find x.
x + (2x – 3) =180
3x – 3 = 180
3x = 177
x = 59
Answer: x = 59
Check : 1 angle is 59, 2nd angle is 115, which leaves 6 degrees for the third

Answer by stanbon(75887)

The diagonal brace on a rectangular gate forms an angle with the horizontal side with degree x and an angle with the vertical side with degree measure 2x – 3. Find x.
x + (2x – 3) =180
3x – 3 = 180
3x = 183
x = 61
Answer: x = 61
Check : 61 + (2*61-3) = 180
61 + 122-3 = 180
183-3=180
180=180
===========
Cheers,
Stan H.